The hospitalization period, in days, for patients following treatment for a certain

type of virus X, where X has the density function f(x) = 4/{π(x2 +1)}

, 0<x <58. Find

the expected value of X. that a person is hospitalized following treatment for this

disorder.


1
Expert's answer
2021-02-01T07:10:29-0500
"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{58}\dfrac{4}{\pi(x^2+1)}dx"

"=\dfrac{4}{\pi}[\arctan(x)]\begin{matrix}
 58 \\
 0
\end{matrix}=\dfrac{4\arctan(58)}{\pi}\not=1"

Therefore


"f(x) = \begin{cases}
 \dfrac{1}{\arctan(58)\cdot(x^2+1)} &\text{if } 0<x<58 \\
 0 &\text{otherwise } 
\end{cases}"

"E(X)=\displaystyle\int_{-\infin}^{\infin}xf(x)dx"


"=\displaystyle\int_{0}^{58}x(\dfrac{1}{\arctan(58)(x^2+1)})dx"

"=\dfrac{1}{2\arctan(58)}[\ln(x^2+1)]\begin{matrix}
 58 \\
 0
\end{matrix}="

"=\dfrac{\ln(3365)}{2\arctan(58)}\approx2.613739"